# Fight like a Woman:

### v2

@@ Conclusion
In this paper, we provide evidence that cases assigned to female judges are significantly more likely to result in conviction than those assigned to male judges with similar career paths. This pattern is, to a large extent, due to the reasons that are specific to domestic violence. The gender difference in conviction rates is stronger for cases involving domestic violence than for crimes that are similar in nature (physical assault) and sentence length (misdemeanors), suggesting the existence of a type of in-group bias as defined by Shayo and Zussman (2011).
Furthermore, we find suggestive evidence for two major drivers of the conviction gender gap in domestic violence. The first is that female judges are stricter than men in cases of intimate partner violence but equally strict for cases of other domestic violence offenses. We claim this is evidence of identification with one’s own gender, as perception of intimate partner violence is highly affected by patriarchal social norms, as opposed to other forms of violence where this might not be the case. This result is consistent with the representational account as defined by Boyd, Epstein, and Martin (2010).
The second driver comes from gender differences in the way judges analyze evidence. There is no gender gap when the defendant is caught red-handed by the police, making the evidence presented clear and objective, but the gap is very strong when there is more room for subjectivity in the analysis. This result is consistent with the informational account as defined by Boyd, Epstein, and Martin (2010).
Finally, we identify an association between a larger gender gap and weaker career pressures for more senior judges in the final stage of the trial court. We believe this to be an important avenue for future research.
@@ Appendix A Additional Details about the Criminal Cases Data
The final dataset was created from two initial datasets.
@@
Using the normalized counts, we train an L1-Regularized Logistic Regression using our training sample. We then validate this algorithm using our three validation samples and find that it correctly classifies 97.9% of the overall cases, 97.9% of the domestic violence cases, and 95.3% of the physical assault cases. Given this high success rate, we use the L1-Regularized Logistic Regression algorithm to define the treatment variable in our full sample.
@@
Having designed the above algorithm, we use it to define the trial judge’s treatment variable $Y$ in the full sample. First, we find which defendants died during their trials and drop them from our sample. We then use the second and third algorithms to define which cases were dismissed and which cases are associated with a non-prosecution agreement. Moreover, we use the trained L1-regularized Logistic Regression algorithm to classify the remaining cases into the categories “defendant is guilty” and “defendant was acquitted”.
Appendix B Final-stage Judges Face Weaker Career Incentives
In Section 6, we find that only final-stage judges present a gender conviction rate gap for domestic violence cases and conclude that weaker career incentives for senior judges are one possible force driving the gender leniency gap. This interpretation relies on the assumption that early-stage judges face stronger career incentives than final-stage judges. The current appendix verifies the plausibility of this assumption.
To do so, we estimate the percentage of cases in which the assigned judge applied for a promotion opportunity within two years after the case assignment. Table B.1 shows these probabilities by crime type and judge characteristics. The first row focuses on the set of cases in which the assigned judge was not in the final stage of the Trial Court System, while the second row focuses on the set of cases in which the assigned judge was in the final stage of the Trial Court System. Odd columns restrict the sample to cases attributed to female judges while even columns restrict the sample to cases attributed to male judges. The header indicates the sample used to estimate these probabilities.
Table B.1: Percentage of Cases in which the Assigned Judge Applied for a Promotion Opportunity within Two Years by Crime Type and Judge Characteristics
Sample: Sample: Sample:
Judge Domestic Violence Physical Assault Misdemeanors
Career Stage Female Male Female Male Female Male
(1) (2) (3) (4) (5) (6)
Early 46.7% 50.9% 46.2% 49.2% 41.4% 48.4%
Final 0.0% 3.9% 0.0% 3.7% 0.0% 2.0%
  • Notes: This table shows the probability of each type of case being assigned to a judge who applied for a promotion opportunity within two years of the day the case was allocated to them. The first row shows this statistic for the set of cases that were attributed to judges who were not in the final stage of the Trial Court System (i.e., substitute judges or magistrates in the initial or intermediate stages). The second row shows this statistic for the set of crimes attributed to judges in the final stage of the Trial Court System. Odd columns restrict the sample to cases attributed to female judges. Even columns restrict the sample to cases attributed to male judges. Columns (1)-(2) refer to cases in the domestic violence sample. Columns (3)-(4) use the sample of physical assault cases. Columns (5)-(6) refer to cases from the misdemeanors sample.
We find that final-stage judges are much less likely to apply for promotion opportunities than early-stage judges. This result suggests that early-stage judges are more eager to be promoted and may face stronger career incentives than final-stage judges. Consequently, the assumption required to interpret weaker career incentives as one of the driving forces of the gender conviction rate gap is plausible.
Lastly, we briefly discuss whether the different conviction criteria adopted by male and female judges may have career consequences for the trial judges. As explained by Laneuville (2024), Appeals Court Judges vote on which trial judge to promote. Consequently, adopting sentencing criteria that lead to more appeal processes and reversed decisions may impact the judge’s career prospects.
To analyze the association between appeal processes, decision reversals, judges’ genders, and judges’ decisions, we regress an outcome variable indicating whether the trial judge’s decision was appealed or reversed on an indicator of a female judge, an indicator of a conviction decision, the interaction of the last two indicators, and district-by-quarter fixed effects. Table B.2 shows the results for the regression whose outcome variable indicates whether the trial judge’s decision was appealed while Table B.3 shows the results for the regression whose outcome variable indicates whether the trial judge’s decision was reversed.
Table B.2: Association between Conviction Sentences and Appealing of Judges’ Decisions on Domestic Violence Cases
Final Early
Appealed Appealed Appealed Appealed
(1) (2) (3) (4)
Fem $0.17^{***}$ $0.09^{***}$ $-0.01$ $0.03$
$(0.03)$ $(0.03)$ $(0.06)$ $(0.05)$
Convicted $0.45^{***}$ $0.41^{***}$
$(0.03)$ $(0.04)$
Fem x Convicted $-0.06$ $-0.09$
$(0.05)$ $(0.07)$
Mean men $0.21$ $0.21$ $0.29$ $0.29$
Quarter-District FE $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Num. obs. $2,599$ $2,599$ $1,749$ $1,749$
${}^{***}p<0.01$; ${}^{**}p<0.05$; ${}^{*}p<0.1$
  • Notes: This table focuses on domestic violence cases and shows the association between a judge’s decision being appealed and the characteristics of the judges and of the decision. Columns (1) and (2) show the results for the sample of cases handled by final-stage judges. Columns (3) and (4) show the results for cases handled by early-stage judges.
Table B.2 shows that female senior judges face appeal processes more frequently than male judges (Column (1)) when analyzing domestic violence cases. It also shows that this higher probability of facing an appeal process is partially explained by convicting more cases (Column (2)). Moreover, it shows that early-stage judges, whose conviction probabilities do not vary by gender according to Table 7, present appeal probabilities that do not vary by gender either (Columns (3) and (4)). Table B.3 presents qualitatively similar results for the probability of having a decision reversed.
Table B.3: Association between Conviction Sentences and Reversal of Judges’ Decisions on Domestic Violence Cases
Final Early
Reversed Reversed Reversed Reversed
(1) (2) (3) (4)
Fem $0.02^{*}$ $0.02^{*}$ $-0.01$ $0.02$
$(0.01)$ $(0.01)$ $(0.06)$ $(0.02)$
Convicted $0.03^{***}$ $0.04^{**}$
$(0.01)$ $(0.01)$
Fem x Convicted $-0.01$ $-0.02$
$(0.02)$ $(0.02)$
Mean men $0.02$ $0.02$ $0.03$ $0.03$
Quarter-District FE $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Num. obs. $2,599$ $2,599$ $1,749$ $1,749$
${}^{***}p<0.01$; ${}^{**}p<0.05$; ${}^{*}p<0.1$
  • Notes: This table focuses on domestic violence cases and shows the association between the reversal of a judge’s decision and the characteristics of the judges and of the decision. Columns (1) and (2) show the results for the sample of cases handled by final-stage judges. Columns (3) and (4) show the results for cases handled by early-stage judges.
We interpret these results as suggestive evidence that the gender conviction rate gap for domestic violence cases may have negative career consequences. For this reason, only female final-stage judges, who face weak career incentives, differ from their male counterparts when analyzing this type of case.
Appendix C Modelling the Gender Conviction-Rate Gap
@@
In this appendix, we propose a simple economic model to formalize the concepts of the gender conviction-rate gap, in-group bias (Shayo and Zussman, 2011) and their driving forces, the representational and informational accounts (Boyd, Epstein, and Martin, 2010). First, to formally define the first two concepts, we adapt the potential outcome model (Rubin, 1974) to a scenario where cases of different types of crime may be treated with a female judge instead of a male judge. Second, to formally define the representational and informational accounts of the in-group bias, we adapt the threshold-crossing model (Heckman and Vytlacil, 2005) to a context where the latent heterogeneity is associated with the evidence of a criminal case and the propensity score is associated with the judge’s conviction criteria.
@@
$\mathbb{E}\left[Y\left(1,1\right)-Y\left(0,1\right)\right]\neq 0,$ (C.1)
@@
$\left\{\begin{matrix}[l]\mathbb{E}\left[Y\left(1,1\right)-Y\left(0,1\right)\right]&>&0\\ \mathbb{E}\left[Y\left(1,1\right)-Y\left(0,1\right)\right]&>&\mathbb{E}\left[Y\left(1,0\right)-Y\left(0,0\right)\right]\end{matrix}\right.,$ (C.2)
@@
$Y=\mathbf{1}\left\{P\left(F,D,W\right)\leq E\left(F,D,R,V\right)\right\},$ (C.3)
@@
Note that Equation (C.3) imposes two exclusion restrictions. While the relationship status between the victim and the defendant — $W$ —- only enters the judge’s conviction criterion function, being caught red-handed — $R$ — only enters the evidence interpretation function. The first exclusion restriction is imposed because the Criminal Process Code does not include this variable as valid evidence against the defendant, but the Criminal Law Code includes this variable as an aggravating factor when setting the sentence. The second exclusion restriction is imposed because the Criminal Process Code considers in flagrante delicto as a piece of evidence against the defendant but not as an aggravating factor.
@@
$P\left(1,1,1\right)-P\left(0,1,1\right)<P\left(1,1,0\right)-P\left(0,1,0\right)\leq 0,$ (C.4)
@@
$E\left(1,1,0,v\right)-E\left(0,1,0,v\right)>E\left(1,1,1,v\right)-E\left(0,1,1,v\right)\geq 0,$ (C.5)
@@
In the main text, we use data from criminal offenses in the State of Sao Paulo, Brazil, to test the existence of each one of the four concepts defined above. In Table 3, we show that the gender conviction-rate gap (Equation (C.1)) exists. Then, in Table 4, we show that in-group bias (Equation (C.2)) is present in our application. Finally, Table 5 shows that our data is consistent with the representational account (Equation (C.4) while Table 6 provides evidence that our data is consistent with the informational account (Equation (C.5).

### v3

@@ Conclusion
In this paper, we provide evidence that domestic violence cases assigned to female judges are significantly more likely to result in conviction than those assigned to male judges with similar career paths. This pattern is, to a large extent, due to the reasons that are specific to domestic violence cases. The gender difference in conviction rates is stronger for cases involving domestic violence than for crimes that are similar in nature (physical assault) and sentence length (misdemeanors), suggesting the existence of a type of in-group bias as defined by Shayo and Zussman (2011).
Furthermore, we find suggestive evidence for two major drivers of the gender conviction-rate gap in domestic violence. The first driver is that female judges are stricter than men in cases of intimate partner violence but equally strict in cases of other types of domestic violence. We claim this finding shows the relevance of identification with one’s own gender, as the perception of intimate partner violence is highly affected by patriarchal social norms, while other forms of violence are not affected by gender stereotypes. This result is consistent with the representational account as defined by Boyd, Epstein, and Martin (2010).
The second driver comes from gender differences in the way judges analyze evidence. There is no gender gap when the defendant is caught red-handed by the police (clear and objective piece of evidence), but the gap is very strong when there is more room for subjectivity in the analysis. This result is consistent with the informational account as defined by Boyd, Epstein, and Martin (2010).
Lastly, we find that the gender conviction-rate gap has no impact on the probability of appeals, ruling reversals or recidivism for female judges in domestic violence cases. Consequently, we conclude that there is no evidence that the gender conviction-rate gap affects the use of Courts’ resources or the defendants’ future criminal behavior.
@@ Appendix A Additional Details about the Criminal Cases Data
The final dataset was created from three initial datasets.
@@
  3. 3.
    CPOSG (“Consulta de Processos de Segundo Grau”): It contains information about all appealing criminal cases in TJ-SP between 2010 and 2019.
@@
Using the normalized counts, we train an L1-Regularized Logistic Regression using our training sample. We then validate this algorithm using our three validation samples and find that it correctly classifies 97.9% of the overall cases, 97.9% of the domestic violence cases, and 95.3% of the physical assault cases. Given this high success rate, we use the L1-Regularized Logistic Regression algorithm to classify which trial judge rulings concluded that the “defendant is guilty” in our full sample.
@@
Having designed the above algorithm, we use it to define whether the trial judge reached a conviction ruling, which is our main outcome variable. First, we find which defendants died during their trials and drop them from our sample. We then use the second and third algorithms to define which cases were dismissed and which cases are associated with a non-prosecution agreement. Moreover, we use the trained L1-regularized Logistic Regression algorithm to classify the remaining cases into the categories “defendant is guilty” and “defendant was acquitted”. In the end, our main outcome value equals $1$ if the case was classified into the category “defendant is guilty” and equals 00 otherwise.
We also use the CPOPG dataset to measure recidivism (Section 6.2). We focus on whether the defendant recidivated within 2 years of their case’s final sentence. Following Possebom (2023), a defendant $i$ in a case $j$ recidivated ($Y_{ij}=1$) if and only if defendant $i$’s full name appears in a case $\bar{j}$ whose starting date is within 2 years after case $j$’s final sentence’s date. Importantly, case $\bar{j}$ can be about any type of crime, including very severe crimes, while case $j$ has to be about domestic violence, misdemeanor, or physical assault offenses. To match defendants’ names across cases, we use the Jaro–Winkler similarity metric and define a match if the similarity between full names in two different cases is greater than or equal to 0.95.181818Abramitzky2019 match full names in historical Censuses in the U.S. and Norway. They define a match between two individuals if the Jaro–Winkler similarity between their names is greater than or equal to 0.90 and if their dates of birth match exactly. Since we do not observe defendants’ dates of birth, we adopt a stricter Jaro-Winkler similarity threshold to define a match in our dataset.
Moreover, in Table 8’s even columns, we exclude from the recidivism definition any case that is a domestic violence case. Specifically, we observe the offense type of case $\bar{j}$ and do not count it as a recidivism event if it is a domestic violence offense.
Importantly, in Section 6.2 only, we delete the case-defendant pairs whose cases started in 2018 and 2019 because their recidivism variable is not properly defined due to right-censoring.
Lastly, we use the CPOSG dataset to measure whether a case was analyzed by the Appeals Court and whether the Appeals Court’s ruling reverses the Trial Judge’s ruling.
We start by merging the CJPG dataset with the CPOSG dataset using each case’s id code. When merging these datasets, we create an indicator variable that denotes which cases were analyzed by the Appeals Court, i.e., which cases were matched. This is the outcome variable in Table 6.1’s odd columns.
We, then, randomly select 50 cases per year for manual classification (2010-2019) and divide them into three categories: “cases that went to the Appeals Court, but were immediately returned due to bureaucratic errors”, “cases whose trial judge’s rulings were affirmed” and “cases whose trial judge’s rulings were reversed”.
Now, we use those 500 manually classified cases to train a classification algorithm. To do so, we divide them into a training sample (300 cases) and a validation sample (200 sentences).
First, we design an algorithm to identify which cases went to the Appeals Court but were immediately returned. To do so, we simply check whether the Appeals Court’s decision is missing in our dataset.
Finally, we design an algorithm to classify the non-empty cases into two categories: “cases whose trial judge’s sentences were affirmed” and “cases whose trial judge’s sentences were reversed”. To do so, we define a bag of words that were found to be strong signals of sentence reversal when we manually classified the cases in our sample. [^19] Possebom (2023, Appendix I.2) lists the entire bag of words. We, then, count how many times each one of those expressions appears in each sentence, and we normalize those counts to be between 0 and 1.
Using the normalized counts, we train an L1-Regularized Logistic Regression using our training sample. We then validate this algorithm using our validation sample and find that it correctly classifies 96.2% of the cases. Given this high success rate, we use the L1-Regularized Logistic Regression algorithm to classify which trial judge’s rulings were reversed in our full sample. This is the outcome variable in Table 6.1’s even columns.
Appendix B Modelling the Gender Conviction-Rate Gap
@@
In this appendix, we propose a simple economic model to formalize the concepts of gender conviction-rate gap and in-group bias (Shayo and Zussman, 2011), and their driving forces, the representational and informational accounts (Boyd, Epstein, and Martin, 2010). First, to formally define the first two concepts, we adapt the potential outcome model (Rubin, 1974) to a scenario where cases of different types of crime may be treated with a female judge instead of a male judge. Second, to formally define the representational and informational accounts of the in-group bias, we adapt the threshold-crossing model (Heckman and Vytlacil, 2005) to a context where the latent heterogeneity is associated with the evidence of a criminal case and the propensity score function is associated with the judge’s conviction criteria.
@@
$\mathbb{E}\left[Y\left(1,1\right)-Y\left(0,1\right)\right]\neq 0,$ (B.1)
@@
$\left\{\begin{matrix}[l]\mathbb{E}\left[Y\left(1,1\right)-Y\left(0,1\right)\right]&>&0\\ \mathbb{E}\left[Y\left(1,1\right)-Y\left(0,1\right)\right]&>&\mathbb{E}\left[Y\left(1,0\right)-Y\left(0,0\right)\right]\end{matrix}\right.,$ (B.2)
@@
$Y=\mathbf{1}\left\{P\left(F,D,W\right)\leq E\left(F,D,R,V\right)\right\},$ (B.3)
@@
Note that Equation (B.3) imposes two exclusion restrictions. While the relationship status between the victim and the defendant — $W$ —- only enters the judge’s conviction criterion function, being caught red-handed — $R$ — only enters the evidence interpretation function. The first exclusion restriction is imposed because the Criminal Process Code does not include this variable as valid evidence against the defendant, but the Criminal Law Code includes this variable as an aggravating factor when setting the sentence. The second exclusion restriction is imposed because the Criminal Process Code considers in flagrante delicto as a piece of evidence against the defendant but not as an aggravating factor.
@@
$P\left(1,1,1\right)-P\left(0,1,1\right)<P\left(1,1,0\right)-P\left(0,1,0\right)\leq 0,$ (B.4)
@@
$E\left(1,1,0,v\right)-E\left(0,1,0,v\right)>E\left(1,1,1,v\right)-E\left(0,1,1,v\right)\geq 0,$ (B.5)
@@
In the main text, we use data from criminal offenses in the State of Sao Paulo, Brazil, to test the existence of each one of the four concepts defined above. In Table 3, we show that the gender conviction-rate gap (Equation (B.1)) exists. Then, in Table 4, we show that in-group bias (Equation (B.2)) is present in our application. Finally, Table 5 shows that our data is consistent with the representational account (Equation (B.4) while Table 6 provides evidence that our data is consistent with the informational account (Equation (B.5).